# Casimir effect with dielectric matter in salted water and implications at the cell scale

### v1

@@ I Introduction
$\left(\frac{1}{2}+\overline{n}_{\omega}\right)\hbar\omega=\frac{\hbar\omega}{2}\coth\frac{\hbar\omega}{2k_{\mathrm{B}}T}\quad\underset{T\to\infty}{\longrightarrow}\;k_{\mathrm{B}}T+O\left(\frac{1}{T}\right)~.$ (1)
@@
$\frac{1}{2}\coth\frac{\hbar\omega}{2k_{\mathrm{B}}T}=\sum_{n=-\infty}^{+\infty}\frac{k_{\mathrm{B}}T}{\hbar\left(\omega-\imath{\xi_{n}}\right)}\quad,\quad\xi_{n}=n\frac{2\pi k_{\mathrm{B}}T}{\hbar}~.$ (2)
@@ II Scattering theory of Casimir interaction in electrolytes
$\mathscr{F}_{0}=\mathscr{F}_{\mathrm{univ}}^{\mathrm{bulk}}+\mathscr{F}_{0}^{(\ell)}\quad,\quad\mathscr{F}_{\mathrm{univ}}^{\mathrm{bulk}}=-\frac{A~H}{12\pi\,d^{2}}\quad,\quad H=\frac{3\,\zeta(3)}{4}\,k_{\mathrm{B}}T\,.$ (6)
@@ IV Universal Casimir interaction in the two-spheres or two-cylinders geometries
$\mathscr{F}_{\mathrm{univ}}^{\mathrm{sph}}=-k_{\mathrm{B}}T\,f\left(d,R_{1},R_{2}\right)~.$ (9)
@@
$x=\frac{d}{R_{\mathrm{eff}}}\quad,\quad u=\frac{R_{1}\,R_{2}}{\left(R_{1}+R_{2}\right)^{2}}~,$ (12)
@@
$y=\frac{\left(d+R_{1}+R_{2}\right)^{2}+R_{1}^{2}+R_{2}^{2}}{2R_{1}\,R_{2}}=1+x\left(1+\frac{ux}{2}\right)~.$ (13)
@@
$\mathscr{F}_{\mathrm{univ}}^{\mathrm{cyl}}=-k_{\mathrm{B}}T\,\frac{L}{d}\,\phi\left(d,R_{1},R_{2}\right)~.$ (14)
@@
In order to discuss the implications of these results for biological systems in section VI, we depict in Fig. 5b the function $\phi$ for two cylinders with equal radii $R_{1}=R_{2}=R$. The function $\phi$ is plotted versus the geometric parameter $x$ defined as in (12), but for cylinders. The grayed out band in the figure shows where $\phi<0.001$, that is where the binding energy $|\mathscr{F}|$ is smaller than the Brownian motion energy scale for L=1000dL=1000~d (see the discussions to be presented in section VI). The take-home message from the figure is that the universal Casimir free energy dominates the Brownian motion energy $k_{\mathrm{B}}T$ imposed by the surrounding water over a broad range of parameters, including in particular distances $d$ of the order of $R_{\rm eff}$. The striking difference between the two-spheres case (Fig. 5a) and the two-cylinders one (Fig. 5b) is simply due to the presence of the large factor $\frac{L}{d}$ in formula (14) which makes $|\mathscr{F}_{\mathrm{univ}}^{\mathrm{cyl}}|$ larger than $k_{\mathrm{B}}T$ in a wider range of parameters.
@@
$\phi_{\mathrm{PFA}}=\frac{\zeta(3)}{32}\,\sqrt{\frac{2R_{\mathrm{eff}}}{d}}\quad,\quad\phi_{\mathrm{SSA}}=\frac{891\pi}{4096}\,\frac{R_{1}^{2}~R_{2}^{2}}{d^{4}}~.$ (15)
@@ V Non-universal contributions with dielectric matter in salted water
In order to get reliable estimations for the non universal contributions to interaction between cylinders, we use the fact that these contributions are expected to be small due to the index-matching at most non null Matsubara frequencies and, therefore, that they should be approximately represented by a pairwise summation technique [3]. We stress that such an approximation cannot give an accurate description of the universal contribution as perturbation theory is not at all adapted to this case. Guided by the pairwise summation, we shall take an educated guess and assume that the order of magnitude of the ratio between universal and non universal contributions arising from two cylinders of radii $R$ can be roughly estimated from the same calculation in a setup of two slabs of thickness $w\equiv 2R$ separated by a gap $d$ of salted water.
@@ VI Implications at the cell scale
We also emphasize that this magnitude has been obtained without any ad hoc fine-tuning of the parameters. The energy constants are given by the theory of the universal Casimir interaction reviewed in this paper, and they do not depend on any specific detail for the biological matter involved. Hence the binding energy depends only on geometrical parameters known for bundles of actin filaments [74]. Note also that the numbers presented for actin (red marker shown on Fig. 9) correspond to $x=4$ in Fig. 5b, where the PFA expression largely overestimates the correct value of $\phi$. This implies that the calculation presented in [66] for cylinders was mandatory for obtaining a reliable estimation for bundles of actin filaments.
Another important point is clear after the discussions in the present paper. We have given an estimation of the non universal Casimir interaction in §V, and shown that it is more than a hundred times smaller than the universal one. It is thus clear that the non universal interaction is not only much smaller than the universal one, but also much too small to have any relevance in a medium dominated by Brownian agitation. It is only the presence of the universal contribution which makes the Casimir binding energy important for biomechanics at the cell scale.

### v2

@@ I Introduction
$\left(\frac{1}{2}+\overline{n}_{\omega}\right)\hbar\omega=\frac{\hbar\omega}{2}\coth\frac{\hbar\omega}{2k_{\mathrm{B}}T}\quad\underset{T\to\infty}{\longrightarrow}\;k_{\mathrm{B}}T+O\left(\frac{1}{T}\right)\penalty 10000\ .$ (1)
@@
$\frac{1}{2}\coth\frac{\hbar\omega}{2k_{\mathrm{B}}T}=\sum_{n=-\infty}^{+\infty}\frac{k_{\mathrm{B}}T}{\hbar\left(\omega-\imath{\xi_{n}}\right)}\quad,\quad\xi_{n}=n\frac{2\pi k_{\mathrm{B}}T}{\hbar}\penalty 10000\ .$ (2)
@@ II Scattering theory of Casimir interaction in electrolytes
$\mathscr{F}_{0}=\mathscr{F}_{\mathrm{univ}}^{\mathrm{bulk}}+\mathscr{F}_{0}^{(\ell)}\quad,\quad\mathscr{F}_{\mathrm{univ}}^{\mathrm{bulk}}=-\frac{A\penalty 10000\ H}{12\pi\,d^{2}}\quad,\quad H=\frac{3\,\zeta(3)}{4}\,k_{\mathrm{B}}T\,.$ (6)
@@ IV Universal Casimir interaction in the two-spheres or two-cylinders geometries
$\mathscr{F}_{\mathrm{univ}}^{\mathrm{sph}}=-k_{\mathrm{B}}T\,f\left(d,R_{1},R_{2}\right)\penalty 10000\ .$ (9)
@@
$x=\frac{d}{R_{\mathrm{eff}}}\quad,\quad u=\frac{R_{1}\,R_{2}}{\left(R_{1}+R_{2}\right)^{2}}\penalty 10000\ ,$ (12)
@@
$y=\frac{\left(d+R_{1}+R_{2}\right)^{2}+R_{1}^{2}+R_{2}^{2}}{2R_{1}\,R_{2}}=1+x\left(1+\frac{ux}{2}\right)\penalty 10000\ .$ (13)
@@
$\mathscr{F}_{\mathrm{univ}}^{\mathrm{cyl}}=-k_{\mathrm{B}}T\,\frac{L}{d}\,\phi\left(d,R_{1},R_{2}\right)\penalty 10000\ .$ (14)
@@
In order to discuss the implications of these results for biological systems in section VI, we depict in Fig. 5b the function $\phi$ for two cylinders with equal radii $R_{1}=R_{2}=R$. The function $\phi$ is plotted versus the geometric parameter $x$ defined as in (12), but for cylinders. The grayed out band in the figure shows where $\phi<0.001$, that is where the binding energy $|\mathscr{F}|$ is smaller than the Brownian motion energy scale for $L=1000\penalty 10000\ d$ (see the discussions to be presented in section VI). The take-home message from the figure is that the universal Casimir free energy dominates the Brownian motion energy $k_{\mathrm{B}}T$ imposed by the surrounding water over a broad range of parameters, including in particular distances $d$ of the order of $R_{\rm eff}$. The striking difference between the two-spheres case (Fig. 5a) and the two-cylinders one (Fig. 5b) is simply due to the presence of the large factor $\frac{L}{d}$ in formula (14) which makes $|\mathscr{F}_{\mathrm{univ}}^{\mathrm{cyl}}|$ larger than $k_{\mathrm{B}}T$ in a wider range of parameters.
@@
$\phi_{\mathrm{PFA}}=\frac{\zeta(3)}{32}\,\sqrt{\frac{2R_{\mathrm{eff}}}{d}}\quad,\quad\phi_{\mathrm{SSA}}=\frac{891\pi}{4096}\,\frac{R_{1}^{2}\penalty 10000\ R_{2}^{2}}{d^{4}}\penalty 10000\ .$ (15)
@@ V Non-universal contributions with dielectric matter in salted water
In order to get reliable estimations for the non universal contributions to interaction between cylinders, we use the fact that these contributions are expected to be small due to the index-matching at most non null Matsubara frequencies and, therefore, that they should be approximately represented by a pairwise summation technique [3]. We stress that such an approximation cannot give an accurate description of the universal contribution as perturbation theory is not at all adapted to this case. Guided by the pairwise summation, we consider that the order of magnitude of the ratio between universal and non universal contributions arising from two cylinders of radii $R$ can be roughly estimated from the same calculation in a setup of two slabs of thickness $w\equiv 2R$ separated by a gap $d$ of salted water.
@@ VI Implications at the cell scale
We also emphasize that this magnitude has been obtained without any ad hoc fine-tuning of the parameters. The energy constants are given by the theory of the universal Casimir interaction reviewed in this paper, and they do not depend on any specific detail for the biological matter involved. Hence the binding energy depends only on geometrical parameters known for bundles of actin filaments [73]. Note also that the numbers presented for actin (red marker shown on Fig. 9) correspond to $x=4$ in Fig. 5b, where the PFA expression overestimates the correct value of $\phi$ by a factor 13. This implies that the calculation presented in [65] for cylinders was mandatory for obtaining a reliable estimation for bundles of actin filaments.
Another important point is clear after the discussions in the present paper. We have given an estimation of the non universal Casimir interaction in §V, and shown that it is more than a hundred times smaller than the universal one. For the relevant distance d=6d=6nm between actin filaments, the non universal contribution is then smaller than $0.05\,k_{\mathrm{B}}T$. It is thus clear that the non universal interaction is not only much smaller than the universal one, but also much too small to have any relevance in a medium dominated by Brownian agitation. It is only the presence of the universal contribution which makes the Casimir binding energy important for biomechanics at the cell scale.